June 26, 2009TLH, NFMCC Friday Meeting1 MICE Updates MICE News Overview ISIS and MICE Schedules DAQ Review Spectrometer Update Aspirational Schedule Terry Hart, NFMCC Friday Meeting, June 26, 2009 Decay Solenoid Target BM1 BM2 Details at www.mice.iit.edu/vc/vc121/vc121.html,www.mice.iit.edu/vc/vc121/vc121.html
June 26, 2009TLH, NFMCC Friday Meeting2 Brief MICE news [Alain Blondel] 1. Decay solenoid: operations will restart in early July. Criteria for acceptance have been specified at the MICE TB on 22Jun09 2. Target: construction according to plan: installation in August still holds. 3. Spectrometer solenoid: as of 18Jun09 target was to start cooling today 4. Running plan A first go has been prepared in discussion between Marco-Linda (MOMs in sept-oct) and AB (see following) WE WILL NEED SHIFTERS (see later) 5. Data flow workshop next week (1st July?) production manager role will be defined (someone will be wanted!) 6. Wanted: volunteer/candidate for VC organization 7. Transnational Access (for EU) is in place Advertised in EUCARD and MICE front page 8. work in US (MRI proposal) and possibly CERN to reinforce RF power effort.
June 26, 2009TLH, NFMCC Friday Meeting3 ISIS Schedule [Yagmur Torun] From machine scheduling meeting (Jun 23): –User run (cycle 2009/01) ends tonight –Record beam current (230 A) achieved this week (Jun 23) –Machine physics Jun 26-28 –Minor maintenance Jun 29-Jul 2 –Run-up Jul 3-5 –Next run (cycle 2009/02) Jul 7-Aug 13 Shielding decouples MICE from ISIS run, OPN will be extended to cover 2nd run cycle
June 26, 2009TLH, NFMCC Friday Meeting4 MICE Schedule [Yagmur Torun] Painting and electrical work in hall to be finished next week Need to clean up before moving lots of air A/C test with full load (DS compressor on) Decay solenoid in operation early July Remove shielding during August shutdown Hook up magnets in DSA Target installation in vault
June 26, 2009TLH, NFMCC Friday Meeting5 DAQ Review [Jean-Sebastien Graulich] One day review forOne day review for MICE Data Acquisition AND Controls –At RAL on June 4 th 2009 4 external Reviewers4 external Reviewers –Pierre Vande Vyvre (Chair); CERN ALICE –Fritz Bartlett; FNAL D0 –Tim Nicholls; RAL T2K –Emilio Radiccioni; INFN Bari TOTEM 1 Internal Reviewer1 Internal Reviewer –John Cobb; Oxford
June 26, 2009TLH, NFMCC Friday Meeting6 DAQ “Bottom Line” [Jean-Sebastien Graulich] The DAQ Review is Done !The DAQ Review is Done ! Very useful comments have been receivedVery useful comments have been received Detector DAQ appears to be in shape and should allow data taking for Step III and beyondDetector DAQ appears to be in shape and should allow data taking for Step III and beyond Unified Controls, Data Flow and Online Database are more recent projects and the committee has noticed it’s not yet matureUnified Controls, Data Flow and Online Database are more recent projects and the committee has noticed it’s not yet mature –Need better integration DAQ/Controls –More thinking on Online Data Buffer Disk Access Serious concern about the man power available on ControlsSerious concern about the man power available on Controls –Call for attention from the management The Control and DAQ Project is not yet over...The Control and DAQ Project is not yet over...
June 26, 2009TLH, NFMCC Friday Meeting7 Spectrometer Solenoid Update [Steve Virostek] An internal leak had developed between the cryostat and the magnet insulating vacuum in early May –Significant effort and some level of disassembly was required to locate and repair the leak –Leak was determined to be at an internal Conflat flange Magnet is repaired and leak tight and has been under vacuum for the past 6 days Insulating vacuum is ready for magnet cool down (today) The required cryogens (LN and LHe), the FNAL LN supply control system and an improved LHe transfer line (and procedure) are all in place Original first magnet cold mass has been modified and is ready for re- installation of the MLI –Additional cryostat penetrations and extra vent line added –Modification of fill line is complete
June 26, 2009TLH, NFMCC Friday Meeting8 Spectrometer Solenoid Upcoming Tasks [Steve Virostek] Complete the LN and LHe cool down of the magnet (approximately a 1 week process) Run the system at operating temperature for ~24 hours to ensure recondensing system is functional Power the magnet and train coils Warm up for shipping to FNAL